# Bialgebras, and Lie monoid actions in Morse and Floer theory, I

### v1

@@ 1. Introduction
Consider a compact Lie group $G$, or more generally a Lie monoid, acting on a smooth manifold $X$ (resp. a symplectic manifold $M$). How does its action interact with the multiplicative structures of Morse theory of $X$ (resp. Floer theory of $M$)? The goal of this paper is to introduce a new ad hoc algebraic framework in order to answer this question.
@@
As for non-equivariant Floer theory, if one wants to allow more general symplectic manifolds and (possibly immersed) Lagrangians, and drop the “nice enough” and “well-composable at the left” assumptions, one must work with curved structures. Curved analogs in our setting should also exist, and should consist in operations for which entries of the multi-index $\underline{l}$ can be zero. We expect a (by no means straightforward) parallel story as in [FOOO09a, FOOO09b], with bounding chains, etc…
@@
The first author plans to prove Conjecture A and B with due details in forthcoming work [Caza]. In Section1.1 and throughout the paper we have included some hints for the Morse and Floer experts that connect our constructions with these conjectures. Except for these hints, we hope that this paper is accessible without prior knowledge in Morse and Floer theory.
@@ 1.1. Idea of the constructions
Figure 2. The moduli space $K^{3}_{2}$. Each of the six chamber is parametrized by two heights. When a height goes to zero, one crosses a wall (interior dashed lines) and moves to an adjascent chamber. When a height goes to infinity, the graph breaks: one reaches an exterior boundary of the hexagon.
This suggests that more generally, there should exist moduli spaces $K^{k}_{l}$ encoding the combinatorics of similar graphs with $k$ inputs and $l$ outputs. These should contain $K_{k}\times K_{l}\times[0,+\infty)$ by glueing an ascending $k$-tree with a descending $l$-tree, with distance between the root vertices given by the $[0,+\infty)$ parameter. When this parameter goes to zero, this family should continue, with the two merging vertices replaced by a “Hopf pattern” (as in $K^{2}_{2}$), and the new born vertices might also collide with other vertices, giving further Hopf patterns as in the two bottom chambers of $K^{3}_{2}$. In this process, graphs become increasingly complicated, and it was not so clear to us how to describe these in full generality, until André Henriques made a beautiful observation to us: “your graphs are obtained by intersecting trees”, i.e. as in Figure 3. This observation seems to have been known to combinatorics experts as well, as we were told by Bruno Vallette and Frédéric Chapoton.
@@
Approach (1) might also work, and perhaps would correspond to [SU11], see Section 1.2. But further subtleties coming from the fact that the diagonals of the associahedra are non associative (see [MS06]) would later lead to complications that we prefer to avoid (not to mention other subtleties of realizing geometrically and coherently the new boundaries).
@@ 1.2. Related work
$A_{\infty}$-algebras are strong homotopy counterparts of associative algebras: these are algebras over the operad $A_{\infty}$, which is Koszul dual to the operad $\mathcal{A}ss$ encoding associative algebras. Bialgebras are “gebras” over a prop $\mathcal{B}iass$. One would wish a similar Koszul dual prop. Koszul duality for props is delicate, and has only been established for properads [Val07]. Unfortunately $\mathcal{B}iass$ is not a properad, since it is not “quadratic”.
@@
In order to construct a chain-level version of $\mathcal{S}ymp$, Bottman introduced a 2-associahedron [Bot19a] in order to turn $\mathcal{S}ymp$ to an $(A_{\infty},2)$-category [BC21] that contains Fukaya categories as 1-morphism categories. He encountered similar issues of polytopes not being “face-coherent” as explained before, and suggested in [Bot] a strategy to resolve it, which partly inspired our work. A more detailed account will appear in [BW]. Frédéric Chapoton informed us that he also considered using similar multi-index polytopes back in 1998.
In [Caz23], the first author outlined how to extend Wehrheim and Woodward’s constructions one dimension down, i.e. as a $(1..4)$-extended TFT. It overcomes issues of moduli spaces having quotient singularities by working equivariantly, following Manolescu and Woodward’s work [MW12]. There, the target is (the completion of) a “partial222in the sense that some compositions are partially defined 2-category” $\mathcal{H}am$: a real analog of the Moore-Tachikawa category [MT12], in which objects are compact Lie groups, and which contains $\mathcal{S}ymp$ as a subcategory of $\mathrm{End}_{\mathcal{H}am}(\{e\},\{e\})$. We expect:
@@ 2.1. Trees, forests and the interior of the $f$-bimultiplihedron
The following two kinds of forests will play a particular role. We say that $\varphi$ is vertical (or trivial) if it has no vertices, i.e. $\underline{k}=(1,\ldots,1)$, i.e. all edges are infinite. We say that $\varphi$ is almost vertical if it is trivalent and has no internal edges, i.e. $\underline{k}$ has only 1 and 2 as coefficients.
@@
  • If $\varphi\simeq\varphi^{\prime}$333in the sense that the sets of leaves, roots, edges and vertices are in bijection, consistently with orders, sources and targets. and under this identification $L\simeq L^{\prime}$, then we declare $(\varphi,L)\sim(\varphi^{\prime},L^{\prime})$.
@@
  • $v(\underline{k})=\left|\underline{k}\right|-n({\underline{k}})$, the number of vertices corresponding to a trivalent tree.
In order to orient this space, we use the following canonical identification. For each pair of consecutive leaves $l,l^{\prime}$ lying on the same tree (i.e. over the same root $r$) corresponds a unique vertex: this is the first vertex where the paths going from $l,l^{\prime}$ to $r$ meet. There are $v(\underline{k})$ such pairs for the forest $U$, ordered in the obvious way, and $v(\underline{l})$ for $D$, ordered likewise. The map
(2.8) $i^{\underline{k}}_{\underline{l}}\colon J^{\underline{k}}_{\underline{l}}\to\mathbb{R}^{v(\underline{k})}\oplus\mathbb{R}^{v(\underline{l})}$
mapping a biforest to all these heights (first the ascending forest, then the descending one), is a homeomorphism. We use it to give $J^{\underline{k}}_{\underline{l}}$ the structure of an oriented vector space. This will be useful for computing orientations and signs.
@@ 2.2. Geometric realizations
We also still denote $h\colon\left|U\right|\to\mathbb{R}$ and $h\colon\langle U\rangle\to\mathbb{R}$ the obvious “vertical” projections. The graph associated to the biforest is defined to be the fibered product
@@
For Morse and Floer theory experts, we describe the domains that can be used in order to form the moduli spaces involved in Conjectures A, B and C, see Figure 7. In Morse theory (Conjecture A), to be precise, we will actually use $\left|U\right|\times_{\mathbb{R}}\langle D\rangle$ rather than $\Gamma(U,D)$. In Floer (open string) theory, we will consider a thickened version $\langle\!\langle D\rangle\!\rangle$ of $\langle D\rangle$, a Riemann surface made of shredded strips:
@@ 2.3. The interior of the $f$-biassociahedron
The group of symmetries of $J^{\underline{k}}_{\underline{l}}$ preserving $\Gamma(U,D)$ as a metric graph444i.e. the graph with lenghts of its edges. is:
(2.22) $G^{\underline{k}}_{\underline{l}}=\begin{cases}\mathbb{R}^{\tilde{a}}&\text{ if }\tilde{b}=0,\\ \mathbb{R}^{\tilde{b}}&\text{ if }\tilde{a}=0,\\ \mathbb{R}&\text{ otherwise.}\end{cases}$
@@
corresponding to the vertices of the i-th nonvertical tree. Likewise in the second case. In the third case it acts by an overall shift along $(1,\ldots,1)$.
Let then the (interior of the) forest biassociahedron
(2.24) ${K}^{\underline{k}}_{\underline{l}}=J^{\underline{k}}_{\underline{l}}/G^{\underline{k}}_{\underline{l}},$
oriented as a quotient of $J^{\underline{k}}_{\underline{l}}$. That is, as oriented vector spaces, $J^{\underline{k}}_{\underline{l}}=G^{\underline{k}}_{\underline{l}}\oplus{K}^{\underline{k}}_{\underline{l}}$.
When one of the multi-indices is vertical, ${K}^{\underline{k}}_{\underline{l}}$ is just a product of associahedrons, up to a sign that we are going to determine. For a multi-index $\underline{k}$, let us introduce:
@@
(2.26) ${K}^{\underline{k}}_{\underline{l}}\simeq(-1)^{\spadesuit_{\tilde{\underline{k}}}}K_{k_{1}}\times\cdots\times K_{k_{a}}.$
Notice that in (2.26) we really mean $\spadesuit_{\tilde{\underline{k}}}$, recall that $\tilde{\underline{k}}$ stands for $\underline{k}$ with the vertical trees removed.
@@
(2.27) ${K}^{\underline{k}}_{\underline{l}}\simeq(-1)^{\spadesuit_{\tilde{\underline{l}}}}K_{l_{1}}\times\cdots\times K_{l_{b}}.$
@@
Observe first that the obvious ${K}^{\underline{k}}_{\underline{l}}\simeq{K}^{\tilde{\underline{k}}}_{\tilde{\underline{l}}}$ is orientation preserving, since it commutes with the identifications $i^{\underline{k}}_{\underline{l}}$. Assume therefore that all vertical trees have been removed, so that $\tilde{\underline{k}}=\underline{k}$. On the one hand, $K_{k_{1}}\times\cdots\times K_{k_{a}}$ is oriented so that
@@
On the other hand, ${K}^{\underline{k}}_{\underline{l}}$ is oriented so that
@@
The dimensions of $J^{\underline{k}}_{\underline{l}}$ and $K^{\underline{k}}_{\underline{l}}$ are given by
@@
(2.31) $\dim K^{\underline{k}}_{\underline{l}}$ $=v({\underline{k}})+v({\underline{l}})-c(\underline{k},\underline{l}),$
@@
In particular,
  • The space $J^{\underline{k}}_{\underline{l}}$ has dimension zero if and only if both ${\underline{k}}$ and ${\underline{l}}$ are vertical.
  • The space $K^{\underline{k}}_{\underline{l}}$ has dimension zero if and only if one of ${\underline{k}}$ and ${\underline{l}}$ is vertical, and the other is almost vertical (i.e. contains only 1 and 2).
@@ 2.4. Partial compactification
Topologically, ${K}^{\underline{k}}_{\underline{l}}$ and $J^{\underline{k}}_{\underline{l}}$ are just vector spaces, what is interesting is their compactification, and the combinatorics of their boundary. We will outline how to compactify these to compact manifolds with generalized corners (as in Joyce [Joy16]) $\overline{K}^{\underline{k}}_{\underline{l}}$ and $\overline{J}^{\underline{k}}_{\underline{l}}$. These will be obtained by adding moduli spaces of “broken biforests” ${K}^{\underline{\underline{k}}}_{\underline{\underline{l}}}$ and $J^{\underline{\underline{k}}}_{\underline{\underline{l}},i}$ of various codimensions, that we will describe below. We expect these spaces to have polytope realizations, as in [CP22], using “shuffle products”.
@@
The important feature of these spaces is that they are “face-coherent”, i.e. their boundary are products of themselves. This is what allows one to write down coherent relations for the maps they will define, leading to the notions in Section 3.
@@
Figure 8. Glueing $\underline{k}^{1}=(1,2,2,1,3)$ on top of $\underline{k}^{0}=(2,3)$ gives $\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}=(3,6)$.

### v2

@@ 1. Introduction
Consider a compact Lie group $G$, or more generally a Lie monoid111by which we mean a closed smooth manifold with an associative product, acting on a smooth manifold $X$ (resp. a symplectic manifold $M$). How does its action interact with the multiplicative structures of Morse theory of $X$ (resp. Floer theory of $M$)? The goal of this paper is to introduce a new ad hoc algebraic framework in order to answer this question.
@@
As for non-equivariant Floer theory, if one wants to allow more general symplectic manifolds and (possibly immersed) Lagrangians, and drop the “nice enough” and “well-composable at the left” assumptions, one must work with curved structures. Curved analogs in our setting should also exist, and should include operations for which entries of the multi-index $\underline{l}$ can be zero. We expect a (by no means straightforward) parallel story as in [FOOO09a, FOOO09b], with bounding chains, etc…
@@
The first author plans to prove Conjecture A and B with due details in forthcoming work [Cazc, Caza]. In Section1.1 and throughout the paper we have included some hints for the Morse and Floer experts that connect our constructions with these conjectures. Except for these hints, we hope that this paper is accessible without prior knowledge in Morse and Floer theory.
@@ 1.1. Idea of the constructions
Figure 2. The moduli space $K^{3}_{2}$. Each of the six chambers is parametrized by two heights. When a height goes to zero, one crosses a wall (interior dashed lines) and moves to an adjascent chamber. When a height goes to infinity, the graph breaks: one reaches an exterior boundary of the hexagon.
This suggests that more generally, there should exist moduli spaces $K^{k}_{l}$ encoding the combinatorics of similar graphs with $k$ inputs and $l$ outputs. These should contain $K_{k}\times K_{l}\times[0,+\infty)$ by gluing an ascending $k$-tree with a descending $l$-tree, with distance between the root vertices given by the $[0,+\infty)$ parameter. When this parameter goes to zero, this family should continue, with the two merging vertices replaced by a “Hopf pattern” (as in $K^{2}_{2}$), and the new born vertices might also collide with other vertices, giving further Hopf patterns as in the two bottom chambers of $K^{3}_{2}$. In this process, graphs become increasingly complicated, and it was not so clear to us how to describe these in full generality, until André Henriques made a beautiful observation to us: “your graphs are obtained by intersecting trees”, i.e. as in Figure 3. This observation seems to have been known to combinatorics experts as well, as we were told by Bruno Vallette and Frédéric Chapoton.
@@
Approach (1) might also work, and perhaps would correspond to [SU11], see Section 1.2. But further subtleties coming from the fact that the diagonals of the associahedra are not associative (see [MS06]) would later lead to complications that we prefer to avoid (not to mention other subtleties of realizing geometrically and coherently the new boundaries).
@@ 1.2. Related work
$A_{\infty}$-algebras are strong homotopy counterparts of associative algebras: these are algebras over the operad $A_{\infty}$, which is the cobar construction of the cooperad Koszul dual to the operad $\mathcal{A}ss$ encoding associative algebras. Bialgebras are “gebras” over a properad $\mathcal{B}iass$ and there exists a Koszul duality theory for properads [Val07]. Unfortunately the most economical presentation of the properad $\mathcal{B}iass$ is not a homogeneous quadratic so it does not fit directly into the Koszul duality for properads. Therefore the homotopical generalisation of associative bialgebras is a notion much more subtle than $A_{\infty}$-algebras and this theory requires further developments.
@@
In order to construct a chain-level version of $\mathcal{S}ymp$, Bottman introduced a 2-associahedron [Bot19a] in order to turn $\mathcal{S}ymp$ to an $(A_{\infty},2)$-category [BC21] that contains Fukaya categories as 1-morphism categories. He encountered similar issues of polytopes not being “face-coherent” as explained before, and suggested in [Bot] a strategy to resolve it, which partly inspired our work. A more detailed account will appear in [Bot24]. Frédéric Chapoton informed us that he also considered using similar multi-index polytopes back in 1998.
In [Caz23], the first author outlined how to extend Wehrheim and Woodward’s constructions one dimension down, i.e. as a $(1..4)$-extended TFT. It overcomes issues of moduli spaces having quotient singularities by working equivariantly, following Manolescu and Woodward’s work [MW12]. There, the target is (the completion of) a “partial333in the sense that some compositions are partially defined 2-category” $\mathcal{H}am$: a real analog of the Moore-Tachikawa category [MT12], in which objects are compact Lie groups, and which contains $\mathcal{S}ymp$ as a subcategory of $\mathrm{End}_{\mathcal{H}am}(\{e\},\{e\})$. We expect:
@@ 2.1. Trees, forests and the interior of the $f$-bimultiplihedron
The following two kinds of forests will play a particular role. We say that $\varphi$ is vertical (or trivial) if it has no vertices, i.e. $\underline{k}=(1,\ldots,1)$, i.e. all edges are infinite. We say that $\varphi$ is almost vertical if it is not vertical, is trivalent and has no internal edges, i.e. $\underline{k}$ has only 1 and 2 as coefficients (and at least one 2).
@@
  • If $\varphi\simeq\varphi^{\prime}$444in the sense that the sets of leaves, roots, edges and vertices are in bijection, consistently with orders, sources and targets. and under this identification $L\simeq L^{\prime}$, then we declare $(\varphi,L)\sim(\varphi^{\prime},L^{\prime})$.
@@
  • $v(\underline{k})=\left|\underline{k}\right|-n({\underline{k}})$ the generic number of vertices (see below).
  • we will denote $\underline{1}_{a}=(1,\ldots,1)$ the multi-index with $a$ entries equal to 1.
The spaces $U^{\underline{k}}$, $D_{\underline{l}}$ and $J^{\underline{k}}_{\underline{l}}$ can be explicitly described, and oriented, as follows. For a multi-index $\underline{k}$, define the ordered set
(2.8) $\mathrm{Vert}(\underline{k})=\left\{1,\ldots,\left|\underline{k}\right|\right\}\setminus\left\{k_{1},k_{1}+k_{2},\ldots,\left|\underline{k}\right|\right\};$
which has cardinality $v(\underline{k})=\left|\underline{k}\right|-n({\underline{k}})$. Observe that if $\varphi$ is a forest of type $\underline{k}$, one has a surjective map $\mathrm{Vert}(\underline{k})\to\mathrm{Vert}(\varphi)$ which is bijective iff $\varphi$ is trivalent. Indeed, the elements of $\mathrm{Vert}(\underline{k})$ correspond to pairs of consecutive leaves $l,l^{\prime}$ lying on the same tree (i.e. over the same root $r$). Map such a pair to the first vertex where the paths going from $l,l^{\prime}$ to $r$ meet.
It is straightforward to check that the map it induces
(2.9) $i^{\underline{k}}_{\underline{l}}\colon J^{\underline{k}}_{\underline{l}}\to\mathbb{R}^{\mathrm{Vert}(\underline{k})\cup\mathrm{Vert}(\underline{l})}\simeq\mathbb{R}^{v(\underline{k})}\oplus\mathbb{R}^{v(\underline{l})},$
mapping a biforest to all its heights (first the ascending forest, then the descending one); is a homeomorphism. We use it to give $J^{\underline{k}}_{\underline{l}}$ the structure of an oriented vector space. This will be useful for computing orientations and signs.
@@ 2.2. Geometric realizations
We also still denote $h\colon\left|U\right|\to\mathbb{R}$ and $h\colon\langle U\rangle\to\mathbb{R}$ the obvious “vertical” projections. The graph associated to the biforest, which we will refer to as the Henriques intersection, is defined to be the fibered product
@@
For Morse and Floer theory experts, we describe the domains that can be used in order to form the moduli spaces involved in Conjectures A, B and C, see Figure 7. In Morse theory (Conjecture A), to be precise, we will actually use $\left|U\right|\times_{\mathbb{R}}\left|D\right|$ rather than $\Gamma(U,D)$. In Floer (open string) theory, we will consider a thickened version $\langle\!\langle D\rangle\!\rangle$ of $\langle D\rangle$, a Riemann surface made of shredded strips:
@@ 2.3. The interior of the $f$-biassociahedron
The group of symmetries of $J^{\underline{k}}_{\underline{l}}$ preserving $\Gamma(U,D)$ as a metric graph555i.e. the graph with lenghts of its edges. is:
(2.23) $G^{\underline{k}}_{\underline{l}}=\begin{cases}\mathbb{R}^{\tilde{a}}&\text{ if }\underline{l}=\underline{1}_{b},\\ \mathbb{R}^{\tilde{b}}&\text{ if }\underline{k}=\underline{1}_{a},\\ \mathbb{R}&\text{ otherwise.}\end{cases}$
@@
corresponding to the vertices of the i-th nonvertical tree. Likewise in the second case. In the third case it acts by an overall shift along $v=(1,\ldots,1)$.
Definition 2.6.
Let then the (interior of the) forest biassociahedron be defined as:
(2.25) ${K}^{\underline{k}}_{\underline{l}}=\begin{cases}J^{\underline{k}}_{\underline{l}}/G^{\underline{k}}_{\underline{l}}&\text{ if }\dim G^{\underline{k}}_{\underline{l}}=1,\\ \emptyset&\text{ otherwise.}\end{cases}$
We will occasionally consider the variation $\widetilde{K}^{\underline{k}}_{\underline{l}}=J^{\underline{k}}_{\underline{l}}/G^{\underline{k}}_{\underline{l}}$, regardless of $\dim G^{\underline{k}}_{\underline{l}}$. We orient both spaces as quotients of $J^{\underline{k}}_{\underline{l}}$. That is, as oriented vector spaces, $J^{\underline{k}}_{\underline{l}}=G^{\underline{k}}_{\underline{l}}\oplus\widetilde{K}^{\underline{k}}_{\underline{l}}$, and when not empty, $J^{\underline{k}}_{\underline{l}}=G^{\underline{k}}_{\underline{l}}\oplus{K}^{\underline{k}}_{\underline{l}}$. In particular, we identify $K^{\underline{k}}_{\underline{l}}$ with the standard orthogonal complement of $G^{\underline{k}}_{\underline{l}}$ in $J^{\underline{k}}_{\underline{l}}\simeq\mathbb{R}^{v(\underline{k})+v(\underline{l})}$.
Remark 2.7.
The variant $\widetilde{K}^{\underline{k}}_{\underline{l}}$ seems more natural to consider, and is convenient for describing the higher codimensional corner strata. However it is redundant: when $\dim G^{\underline{k}}_{\underline{l}}\geq 2$, $\widetilde{K}^{\underline{k}}_{\underline{l}}$ is a product of associahedrons up to a sign (Proposition 2.8 below). It turns out to be more convenient to set ${K}^{\underline{k}}_{\underline{l}}=\emptyset$ and the associated maps ${\alpha}^{\underline{k}}_{\underline{l}}=0$ in the definition of $f$-bialgebras, and this causes no loss of information.
For a multi-index $\underline{k}$, let us introduce:
@@
(2.27) $\widetilde{K}^{\underline{k}}_{\underline{l}}\simeq(-1)^{\spadesuit_{\tilde{\underline{k}}}}K_{k_{1}}\times\cdots\times K_{k_{a}}.$
Notice the tilde in $\spadesuit_{\tilde{\underline{k}}}$, with $\tilde{\underline{k}}$ standing for $\underline{k}$ with the vertical trees removed.
@@
(2.28) $\widetilde{K}^{\underline{k}}_{\underline{l}}\simeq(-1)^{\spadesuit_{\tilde{\underline{l}}}}K_{l_{1}}\times\cdots\times K_{l_{b}}.$
@@
Observe first that the obvious $\widetilde{K}^{\underline{k}}_{\underline{l}}\simeq\widetilde{K}^{\tilde{\underline{k}}}_{\tilde{\underline{l}}}$ is orientation preserving, since it commutes with the identifications $i^{\underline{k}}_{\underline{l}}$. Assume therefore that all vertical trees have been removed, so that $\tilde{\underline{k}}=\underline{k}$. On the one hand, $K_{k_{1}}\times\cdots\times K_{k_{a}}$ is oriented so that
@@
On the other hand, $\widetilde{K}^{\underline{k}}_{\underline{l}}$ is oriented so that
@@
The dimensions of $J^{\underline{k}}_{\underline{l}}$, $K^{\underline{k}}_{\underline{l}}$ and $\widetilde{K}^{\underline{k}}_{\underline{l}}$ are given by
@@
(2.32) $\dim K^{\underline{k}}_{\underline{l}}$ $=v({\underline{k}})+v({\underline{l}})-1,$
(2.33) $\dim\widetilde{K}^{\underline{k}}_{\underline{l}}$ $=v({\underline{k}})+v({\underline{l}})-c(\underline{k},\underline{l}),$
@@ 2.4. Partial compactification
Topologically, ${K}^{\underline{k}}_{\underline{l}}$ and $J^{\underline{k}}_{\underline{l}}$ are just vector spaces, what is interesting is their compactification, and the combinatorics of their boundary. They can be compactified to compact manifolds with generalized corners (as in Joyce [Joy16]) $\overline{K}^{\underline{k}}_{\underline{l}}$ and $\overline{J}^{\underline{k}}_{\underline{l}}$, by gluing to them moduli spaces of “broken biforests” of various codimensions. We expect these spaces to have polytope realizations, as in [CP22], using “shuffle products”.
@@
The important feature of these spaces is that they are “face-coherent”, i.e. their boundaries are products of themselves. This is what allows one to write down coherent relations for the maps they will define, leading to the notions in Section 3.
@@
Figure 8. Gluing $\underline{k}^{1}=(1,2,2,1,3)$ on top of $\underline{k}^{0}=(2,3)$ gives $\underline{k}=\underline{k}^{1}\sharp\underline{k}^{0}=(3,6)$.